# A Grassroots Architecture to Supplant Global Digital Platforms by a Global Digital Democracy

### v3

@@ Abstract
We present an architectural alternative to global digital platforms termed grassroots, designed to serve the social, economic, civic, and political needs of local digital communities, as well as their federation. Grassroots platforms may offer local communities an alternative to global digital platforms while operating solely on the smartphones of their members, forsaking any global resources other than the network itself. Such communities may form digital economies without initial capital or external credit, exercise sovereign democratic governance, and federate, ultimately resulting in the grassroots formation of a global digital democracy.
1 Introduction
@@
Here we present an architectural alternative to global digital platforms termed grassroots [45, 43]. It is designed to support grassroots platforms that serve the social, economic, civic, and political needs of local digital communities, as well as provide for their federation. Grassroots platforms may offer local alternatives to global digital platforms while operating solely on the smartphones of their members, forsaking any global resources other than the network itself. The grassroots architecture presented here is a major revision and expansion of an earlier proposal with a similar goal [48].
The grassroots architecture incorporates a grassroots protocol stack and employs the blocklace—a generalisation of the blockchain that functions as a universal, Byzantine fault-tolerant, Conflict-free Replicated Data Type (CRDT) [4]. The grassroots protocol stack provides for dissemination [45], equivocation exclusion [24, 29], ordering consensus [24, 31], State Machine Replication/digital social contracts [12], democratic governance [48], and federation [21]. Envisioned grassroots platforms include grassroots social networks [44], grassroots currencies [46], and a grassroots digital economy comprising people as well as legal persons (banks, corporations, cooperatives, communities, municipalities, and states) employing off-chain and on-chain governance. Together, the grassroots protocol stack and platforms aim to provide scalable foundations for grassroots digital communities that emerge locally and federate globally [21]. Such communities may form digital economies without initial capital or external credit, exercise sovereign democratic governance, and federate, ultimately resulting in the grassroots formation of a global digital democracy [48, 21].
In this paper we review the core concepts of the architecture—digital sovereignty, grassroots systems (Figure 1, top black box) and the blocklace (ibid., bottom blue boxes)—and the grassroots protocol stack (ibid., blue boxes), and only reference the grassroots platforms (ibid., brown boxes).
@@ 3.6 Grassroots Federated Assemblies for Global Democratic Governance
An effort to address this problem employs grassroots federation [21]. A federation may include people and/or child federations forms voluntarily and is governed by a small (say 100-member) assembly chosen by sortition from its individual members and the members of the assemblies of its child federations. The assembly of a federation is sovereign to decide on the admission and removal of its members, as well as on the application to join higher-level federations. The strata to be represented by the assembly of a federation are determined by the evolving structure of the federation and sizes of its member assemblies, which may intersect and therefore have to be weighted.
This approach addresses sybils by having the federation built from small local communities in which members know each other to be genuine, and from communities that federate only if they trust each other to be genuine, having sufficient intersection or relationships among them to base this trust on. Large-scale online voting is eschewed as every federation, no matter how large, is governed by an assembly that engages in small-scale democracy, where votes should be taken by a show of hands for transparency, allowing people to see how their representatives have voted.

### v5

@@ Abstract
We present an architectural alternative to global digital platforms termed grassroots, designed to serve the social, economic, civic, and political needs of local digital communities, including their federation. Grassroots platforms may offer local communities an alternative to global digital platforms while operating solely on the smartphones of their members, forsaking any global resources other than the network itself. Such communities may form digital economies without initial capital or external credit, exercise sovereign democratic governance, and federate, ultimately resulting in the grassroots formation of a global digital democracy.
Keywords:
Grassroots Platforms Distributed Systems Blockchain Blocklace Social Networks Cryptocurrencies Consensus Smart contracts Federation Digital Democracy1 Introduction
@@
Here we present an architectural alternative to global digital platforms termed grassroots [50, 48]. It is designed to serve the social, economic, civic, and political needs of local digital communities, as well as provide for their federation. The architecture support grassroots platforms that may offer local alternatives to global digital platforms while operating solely on the smartphones of their members, forsaking any global resources other than the network itself. The grassroots architecture presented here is a major revision and expansion of an earlier proposal with a similar goal [54].
The grassroots architecture incorporates a grassroots protocol stack and employs the blocklace—a generalisation of the blockchain that functions as a universal, Byzantine fault-tolerant, Conflict-free Replicated Data Type (CRDT) [4]. The grassroots protocol stack provides for dissemination [50], equivocation exclusion [25, 30], ordering consensus [25, 32], State Machine Replication/digital social contracts [12], democratic governance [54], and federation [21, 22]. Envisioned grassroots platforms include grassroots social networks [49], grassroots currencies [51], and a grassroots digital economy comprising people as well as legal persons (banks, corporations, cooperatives, communities, municipalities, and states) employing off-chain and on-chain governance. Together, the grassroots protocol stack and platforms aim to provide scalable foundations for grassroots digital communities that emerge locally and federate globally [22]. Such communities may form digital economies without initial capital or external credit, exercise sovereign democratic governance, and federate, ultimately resulting in the grassroots formation of a global digital democracy [54, 21, 22].
In this paper we review the core concepts of the architecture—digital sovereignty, grassroots systems (Figure 1, top black box), mention the desiderata on grassroots programming languages (ibid., green box), explain the blocklace and the grassroots protocol stack (ibid., blue boxes), and only reference the grassroots platforms (ibid., brown boxes).
@@ 2 Architecture Core Concepts
The grassroots architecture is designed to be implemented on people’s smartphones, with server-farms up in the clouds only playing an auxiliary role, as needed. In that regard it is worth noting that today’s smartphones have thousands of times more of the computing power, memory capacity and networking speed of the Unix workstations that were the workhorses of the Internet and the web at their inception. The reasons that today’s role of smartphones on the Internet is only to access the platforms that operate on it, not to run them, are historical, economic, political, and architectural, not their performance. The grassroots architecture aims to remove these barriers through the notion of grassroots platforms.
@@ 2.2 Grassroots Systems
2.3 Grassroots Programming Languages
A programming language for grassroots systems [52] should serve two goals: First, to ease the implementation of grassroots protocols, grassroots applications and grassroots platforms, by providing an abstraction of the underlying machinery they necessitate. Second, to ease the proof that the resulting system is indeed grassroots. To do so, a grassroots programming language must gave at least four characteristics:
1.
Be a concurrent programming language, supporting the dynamic creation of processes, their communication and synchronisation.
2.
Its processes must be independent of any global/shared/third party resource for their execution.
3.
Processes can reconfigure their communication network dynamically. For this, communication channels should be ‘first class objects’, so that one process can send another a message that contains a communication channel with a third process, and the recipient can then communicate with that third process.
4.
Programs should also be ‘first class objects’, so that one process can send a message that contains a program to another, which can then execute the program upon receipt.
In addition, ideally a grassroots programming language should support metaprogramming [60, 46, 42, 34], to ease the implementation of a grassroots operating system and program development environment for the grassroots programming language within the language. We are presently exploring addressing this challenge with a single-reader/single-writer subset of Concurrent Prolog [52, 46, 45].
@@ 3.6 Grassroots Federated Assemblies for Global Democratic Governance
An effort to address this problem employs grassroots federation [21, 22]. A federation may include people and/or child federations forms voluntarily and is governed by a small (say 100-member) assembly chosen by sortition from its individual members and the members of the assemblies of its child federations. The assembly of a federation is sovereign to decide on the admission and removal of its members, as well as on the application to join higher-level federations. The strata to be represented by the assembly of a federation are determined by the evolving structure of the federation and sizes of its member assemblies, which may intersect and therefore have to be weighted.
This approach addresses sybils by having the federation built from small local communities in which members know each other to be genuine, and from communities that federate only if they trust each other to be genuine, for example by having sufficient intersection or actual relationships among them to base this trust on. Large-scale online voting is eschewed as every federation, no matter how large, is governed by an assembly that engages in small-scale democracy, where votes should be taken by a show of hands for transparency, allowing people to see how their representatives have voted.